If f(x)=2x^2 & g(x)=3x^4, which is equivalent form of f(x)*g(x)?
A.6x^8
B.5x^6
C.6x^6
D.5x^8

Answer by jim_thompson5910(35256)

f(x)*g(x) = (2x^2)*(3x^4) = (2*3)*(x^2*x^4) = 6x^(2+4) = 6x^6

So f(x)*g(x) = 6x^6